# Complexity and curvature of pairs of Burch modules and ideals

### v1

@@ Abstract.
The complexity and curvature of a module were first introduced by Avramov to distinguish modules of infinite homological dimension. Later, Avramov–Buchweitz extended the notion of complexity from a single module to that of pairs of modules, which measures the polynomial growth rate of the minimal number of generators of their Ext modules. Dao studied a similar notion of Tor-complexity. Recently, Dey–Ghosh–Saha initiated the study of Ext and Tor curvature of a pair of modules, which measure the exponential growth rates of the corresponding Ext and Tor, respectively. On the other hand, the concept of Burch ideals was introduced by Dao–Kobayashi–Takahashi, motivated by classical work of Burch. This class includes several large and well-studied families of ideals in a Noetherian local ring $(R,\mathfrak{m},k)$. For example, every nonzero ideal of the form $\mathfrak{a}\mathfrak{m}$ (e.g., $\mathfrak{m}^{n}$ for $n\geqslant 1$), and under mild conditions every integrally closed ideal $I$ with $\operatorname{depth}(R/I)=0$, is a Burch ideal. Suppose $I$ and $J$ are Burch ideals such that $I$ is $\mathfrak{m}$-primary. One of the main results of this article is $\operatorname{cx}_{R}(I,J)=\operatorname{tcx}_{R}(I,J)=\operatorname{cx}_{R}(k)$. Moreover, we show that $R$ is complete intersection if and only if $\operatorname{cx}_{R}(I,J)$ or $\operatorname{tcx}_{R}(I,J)$ is finite if and only if $\operatorname{curv}_{R}(I,J)$ or $\operatorname{tcurv}_{R}(I,J)$ is at most $1$. We deduce these results from the corresponding more general results on Burch modules (in the sense of Dey-Kobayashi).
@@ 1. Introduction
The notions of complexity and curvature of a module were originally introduced by Avramov in [Avr89a], [Avr89b] and [Avr96], see Definitions 2.1 and 2.3. These invariants distinguish modules of infinite projective and injective dimensions. Later, in [AB00], Avramov and Buchweitz extended the notion of complexity to pairs of modules $(M,N)$, which measures the polynomial growth rate of minimal number of generators $\mu(\operatorname{Ext}^{n}_{R}(M,N))$ of the Ext modules as $n$ increases. Dao introduced the analogous concept of Tor-complexity in [Dao07]. This measures the polynomial growth rate of $\mu(\operatorname{Tor}_{n}^{R}(M,N))$. Influenced by these, in [DGS24], Dey, Ghosh and Saha developed the concept of Ext and Tor curvatures for a pair of modules; these measure the exponential growth rate of Ext and Tor modules, respectively, see Definition 2.4. Additionally, in [DGS24, Thms. A and B, Cors. 5.5 and 6.6], Dey, Ghosh and Saha established various inequalities between (Ext and Tor) complexity and curvature of pairs of certain CM (Cohen-Macaulay) modules and that of a single module. Along with other applications, they also provided a number of criteria for a local ring to be complete intersection (including hypersurface and regular local ring) via the complexity and curvature of pairs of CM modules satisfying certain inequalities involving multiplicity, see [DGS24, Sec. 7]. In the present article, we study the (Ext and Tor) complexity and curvature of pairs of Burch modules, and establish various (in)equalities. Along the way, we obtain a number of new characterizations of complete intersection local rings (including hypersurface rings). Recall that $R$ is said to be a complete intersection of codimension $c$ if its completion $\widehat{R}=Q/({\bf f})$, where $(Q,\mathfrak{n})$ is a regular local ring, and ${\bf f}\in\mathfrak{n}^{2}$ is a $Q$-regular sequence of length $c$. If $R$ is complete intersection of codimension at most $1$, then $R$ is called a hypersurface.
The class of Burch ideals and modules is an important class of ideals and modules studied in commutative algebra. In [DKT20], Dao, Kobayashi and Takahashi introduced the notion of Burch ideals motivated by some striking results of Burch [burch, p. 946, Thm. 5.(ii), and p. 947, Cor. 1.(ii)]. Later, Dey and Kobayashi extended this concept, and introduced the notion of Burch submodules in [DK23]. An $R$-module $M$ is said to be a Burch submodule of an $R$-module $L$ if $\mathfrak{m}(M:_{L}\mathfrak{m})\neq\mathfrak{m}M$. A Burch submodule of $R$ is called a Burch ideal. An $R$-module $M$ is called Burch if it is isomorphic to a Burch submodule of some $R$-module. There are several well-studied classes of examples of Burch (sub)modules and ideals. Here we list some interesting classes.
@@
    For a submodule $M$ of an $R$-module $L$, if $\mathfrak{m}M\neq 0$, then $\mathfrak{m}M$ is a Burch submodule of $L$, see [DK23, Ex. 3.4]. Moreover, a non-zero ideal $I$ of the form $\mathfrak{a}\mathfrak{m}$ (for some ideal $\mathfrak{a}$) is a Burch ideal. In particular, if $\mathfrak{m}^{n}\neq 0$ for some $n\geqslant 1$, then $\mathfrak{m}^{n}$ is an $\mathfrak{m}$-primary Burch ideal.
@@
    It follows from (1) and [DK23, Lem. 5.2] that Ulrich modules of dimension 1 are Burch.
@@
    If $M$ is Burch, then so is $M\oplus X$ for any $R$-module $X$.
@@
    Under some mild conditions on $R$, it is shown in [GS24, Prop. 4.6] that every integrally closed ideal $I$ such that $\operatorname{depth}(R/I)=0$ is Burch; in particular, $\mathfrak{m}$-primary integrally closed ideals are Burch ideals.
@@
    If $R$ is of prime characteristic $p$, $F$-finite, and $\operatorname{depth}(R)=1$, then the $e$-th Frobenius pushforward ${{}^{e}\!}{R}$ is Burch for all $e\gg 0$, see [CDKMS, 2.4.(vi)]. Recall that ${{}^{e}\!}{R}$ denotes the $R$-module whose underlying abelian group is $R$, and the $R$-action on ${{}^{e}\!}{R}$ is given by $r\cdot x=r^{p^{e}}x$ for $r\in R$ and $x\in{{}^{e}\!}{R}$.
Burch ideals and modules possess several nice homological properties. For instance, they are strong test modules that detect finite homological dimensions via finite vanishing of Ext or Tor ([burch, p. 946, Thm. 5.(ii)]; [DK23, Prop. 3.16]). They characterize regular, complete intersection, Gorenstein, and CM local rings in terms of the corresponding homological dimensions ([burch, p. 947, Cor. 1.(ii)]; [DK23, Cor. 3.19]; [GS24, Thms. 1.1 and 5.1.(6)]). They also satisfy the Huneke–Wiegand conjecture over Gorenstein rings [CDKMS, Thm. 1.2].
Towards understanding growth in minimal free and injective resolutions of Burch modules, it follows from a result of Avramov [Avr96, Thm. 4] that a Burch module (and hence a Burch ideal) has maximal (injective) complexity and curvature; see Proposition 2.20. As a consequence of this and [Avr96, Thm. 3], it follows that $R$ is a complete intersection if (injective) complexity of $M$ is finite, or (injective) curvature of $M$ is at most $1$, where $M$ is some Burch $R$-module. Since $\mathfrak{m}$ is always Burch (unless $R$ is a field), and $\operatorname{cx}_{R}(M)=\operatorname{tcx}_{R}(\mathfrak{m},M)$, $\operatorname{curv}_{R}(M)=\operatorname{tcurv}_{R}(\mathfrak{m},M)$, $\operatorname{inj\,cx}_{R}(M)=\operatorname{cx}_{R}(\mathfrak{m},M)$, $\operatorname{inj\,curv}_{R}(M)=\operatorname{curv}_{R}(\mathfrak{m},M)$, it is natural to ask the following questions.
@@ 2. Preliminaries
[DGS24, Lem. 2.7 and Rmk. 2.2.(1)] For sequences $\{x_{n}\}$ and $\{y_{n}\}$ of non-negative integers,
@@
The complexity of a module was introduced by Avramov in [Avr89a, Def. 3.1] and [Avr89b, Def. 1.1], while the injective complexity, referred to as plexity, was defined in [Avr89b, Def. 5.1]. The study of (injective) curvature was initiated in [Avr96, Sec. 1].
@@
    [AB00, p. 286] The complexity (or Ext-complexity) of $M$ and $N$ is defined to be cx_R(M,N):=cx{μ_R(Ext_R^n(M,N))}_n⩾0.
@@
    [DGS24, Def. 3.2] The curvature (or Ext-curvature) of $M$ and $N$ is defined by curv_R(M,N):=curv{μ_R(Ext_R^n(M,N))}_n ⩾0.
@@
    [DGS24, Def. 3.3] The Tor-curvature of $M$ and $N$ is defined by tcurv_R(M,N):=curv{μ_R(Tor^R_n(M,N))}_n ⩾0.
@@
The Tor-curvature is also defined in [Sega, 3.8] in terms of length of Tor modules. However, the definitions in terms of $\mu$ and in terms of $\lambda$ are equivalent when $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$; see [DGS24, Lem. 3.4].
@@
    [Avr96, Prop. 2] For any $R$-module $M$, the following hold.
@@
    [Avr96, Thm. 3] The following conditions are equivalent:
@@
In view of [Mat89, p. 140, Lem. 2], $\operatorname{Ext}_{R}^{n}(M,k)\cong\operatorname{Ext}_{R/xR}^{n}(M/xM,k)$ and $\operatorname{Ext}_{R}^{n+1}(k,M)\cong\operatorname{Ext}_{R/xR}^{n}(k,M/xM)$ for all $n\geqslant 0$. So $\beta_{n}^{R}(M)=\beta_{n}^{R/xR}(M/xM)$ and $\mu^{n+1}_{R}(M)=\mu^{n}_{R/xR}(M/xM)$ for all $n\geqslant 0$. Hence the desired equalities follow. ∎
@@
(1) It is shown in [AB00, Thm. II] that $\operatorname{cx}_{R}(M)=\operatorname{inj\,cx}_{R}(M)$ and $\operatorname{cx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}$. By Proposition 2.7.(1), $\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(k)<\infty$. Thus $\operatorname{cx}_{R}(M,N)<\infty$. Hence, in view of Remark 2.6.(3), $\operatorname{curv}_{R}(M,N)\leqslant 1$.
(2) The inequalities for (Tor) complexities are shown in [Dao07, Cor. 5.7] and [DGS24, Prop. 7.9.(1)]. Hence, by Remark 2.6.(3), $\operatorname{tcurv}_{R}(M,N)\leqslant 1$. ∎
@@
(1) Consider the corresponding long exact sequence of Tor modules. Hence, the first claim follows from 2.11, while the desired inequalities can be obtained from 2.11, [DGS24, Lem. 3.4] and Lemma 2.2 by computing the lengths of the Tor modules.
@@
    [DK23, Def. 3.1] A submodule $M$ of an $R$-module $L$ is said to be Burch if
@@
    [DKT20, Def. 2.1] A Burch submodule of $R$ is called a Burch ideal of $R$.
@@
[DK23, Lem. 3.6] Let $M$ be a Burch submodule of an $R$-module $L$ such that $\operatorname{depth}(M)>0$. Then there exists an $M$-regular element $x\in\mathfrak{m}$ such that $k$ is a direct summand of $M/xM$.
@@
[Avr96, Thm. 4], [GS24, Thm. 5.1] Let $M$ be a Burch module. Then
@@
[GS24, Lem. 4.8] Let $S=R[X]_{\langle\mathfrak{m},X\rangle}$. Let $M$ be a submodule of an $R$-module $L$. Set $L^{\prime}:=L[X]_{\langle\mathfrak{m},X\rangle}$ and $M^{\prime}:=(M+XL[X])_{\langle\mathfrak{m},X\rangle}$. Then
@@
In the common hypothesis of all the statements in [GS24, Lem. 4.8], it is assumed that $M$ is a Burch submodule of $L$. Note that the Burch property of $M$ is required for the proof of statement (2) in [GS24, Lem. 4.8], and is not needed elsewhere.
@@
(2) Similarly, as in (1), the desired equalities on injective complexity and curvature can be obtained by using [DGS24, Lem. 3.11.(2)] and Lemma 2.8.(2). ∎
@@
2.26.
Let $\mathcal{A}$ be the category of $R$-modules of Gorenstein dimension zero, and $\mathcal{B}$ be the subcategory of projective $R$-modules. Following the notation as in [Aus-buch, p. 9], $\hat{\mathcal{A}}$ and $\hat{\mathcal{B}}$ denote the category of $R$-modules of finite Gorenstein dimension and of finite projective dimension, respectively. Then, in view of [Aus-buch, Thm. 1.1], for each $U$ in $\hat{\mathcal{A}}$, there is an exact sequence $0\rightarrow U\rightarrow V\rightarrow W\rightarrow 0$ with $V$ in $\hat{\mathcal{B}}$ and $W$ in $\mathcal{A}$.

### v2

@@ Abstract.
The complexity and curvature of a module were first introduced by Avramov to distinguish modules of infinite homological dimension. Later, Avramov–Buchweitz extended the notion of complexity from a single module to that of pairs of modules, which measures the polynomial growth rate of the minimal number of generators of their Ext modules. Dao studied a similar notion of Tor-complexity. Recently, Dey–Ghosh–Saha initiated the study of Ext and Tor curvature of a pair of modules, which measure the exponential growth rates of the corresponding Ext and Tor, respectively. On the other hand, the concept of Burch ideals was introduced by Dao–Kobayashi–Takahashi, motivated by the classical work of Burch, and subsequently extended to modules by Dey–Kobayashi. This class includes several large and well-studied families of modules and ideals over a Noetherian local ring $(R,\mathfrak{m},k)$. For example, these include the residue field $k$ as an $R$-module, every nonzero module of the form $\mathfrak{m}M$ (e.g., $\mathfrak{m}^{n}$ for $n\geqslant 1$), and under mild conditions every integrally closed ideal $I$ with $\operatorname{depth}(R/I)=0$. Suppose $I$ and $J$ are Burch ideals such that $I$ is $\mathfrak{m}$-primary. Motivated by Avramov’s result that Burch modules exhibit extremal complexity and curvature, we establish in this article that $\operatorname{cx}_{R}(I,J)=\operatorname{tcx}_{R}(I,J)=\operatorname{cx}_{R}(k)$. Moreover, we show that $R$ is complete intersection if and only if $\operatorname{cx}_{R}(I,J)$ or $\operatorname{tcx}_{R}(I,J)$ is finite if and only if $\operatorname{curv}_{R}(I,J)$ or $\operatorname{tcurv}_{R}(I,J)$ is at most $1$. We deduce these results from the corresponding more general results on Burch modules.
@@ 1. Introduction
The notions of complexity and curvature of a module were originally introduced by Avramov in [4], [5] and [6], see Definitions 2.1 and 2.3. These invariants distinguish modules of infinite projective and injective dimensions. Later, in [3], Avramov and Buchweitz extended the notion of complexity to pairs of modules $(M,N)$, which measures the polynomial growth rate of minimal number of generators $\mu(\operatorname{Ext}^{n}_{R}(M,N))$ of the Ext modules as $n$ increases. Dao introduced the analogous concept of Tor-complexity in [11]. This measures the polynomial growth rate of $\mu(\operatorname{Tor}_{n}^{R}(M,N))$. Influenced by these, in [12], Dey, Ghosh and Saha developed the concept of Ext and Tor curvatures for a pair of modules; these measure the exponential growth rate of Ext and Tor modules, respectively, see Definition 2.4. Additionally, in [12, Thms. A and B, Cors. 5.5 and 6.6], Dey, Ghosh and Saha established various inequalities between (Ext and Tor) complexity and curvature of pairs of certain CM (Cohen-Macaulay) modules and that of a single module. Along with other applications, they also provided a number of criteria for a local ring to be complete intersection (including hypersurface and regular local ring) via the complexity and curvature of pairs of CM modules satisfying certain inequalities involving multiplicity, see [12, Sec. 7]. In the present article, we study the (Ext and Tor) complexity and curvature of pairs of Burch modules, and establish various (in)equalities. Along the way, we obtain a number of new characterizations of complete intersection local rings (including hypersurface rings). Recall that $R$ is said to be a complete intersection of codimension $c$ if its completion $\widehat{R}=Q/({\bf f})$, where $(Q,\mathfrak{n})$ is a regular local ring, and ${\bf f}\in\mathfrak{n}^{2}$ is a $Q$-regular sequence of length $c$. If $R$ is complete intersection of codimension at most $1$, then $R$ is called a hypersurface.
The class of Burch ideals and modules is an important class of ideals and modules studied in commutative algebra. In [10], Dao, Kobayashi and Takahashi introduced the notion of Burch ideals motivated by some striking results of Burch [7, p. 946, Thm. 5.(ii), and p. 947, Cor. 1.(ii)]. Later, Dey and Kobayashi extended this concept, and introduced the notion of Burch submodules in [13]. An $R$-module $M$ is said to be a Burch submodule of an $R$-module $L$ if $\mathfrak{m}(M:_{L}\mathfrak{m})\neq\mathfrak{m}M$. A Burch submodule of $R$ is called a Burch ideal. An $R$-module $M$ is called Burch if it is isomorphic to a Burch submodule of some $R$-module. There are several well-studied classes of examples of Burch (sub)modules and ideals. Here we list some interesting classes.
@@
    For a submodule $M$ of an $R$-module $L$, if $\mathfrak{m}M\neq 0$, then $\mathfrak{m}M$ is a Burch submodule of $L$, see [13, Ex. 3.4]. Moreover, a non-zero ideal $I$ of the form $\mathfrak{a}\mathfrak{m}$ (for some ideal $\mathfrak{a}$) is a Burch ideal. In particular, if $\mathfrak{m}^{n}\neq 0$ for some $n\geqslant 1$, then $\mathfrak{m}^{n}$ is an $\mathfrak{m}$-primary Burch ideal.
@@
    It follows from (1) and [13, Lem. 5.2] that Ulrich modules of dimension 1 are Burch.
@@
    For two $R$-modules $M$ and $N$, the direct sum $M\oplus N$ is Burch if and only if at least one of $M$ or $N$ is Burch. Since $k^{\oplus\mu(\mathfrak{m})}\cong\mathfrak{m}/\mathfrak{m}^{2}=\mathfrak{m}(R/\mathfrak{m}^{2})$, in view of (1), we conclude that $k$ is a Burch $R$-module.
@@
    Under some mild conditions on $R$, it is shown in [14, Prop. 4.6] that every integrally closed ideal $I$ such that $\operatorname{depth}(R/I)=0$ is Burch; in particular, $\mathfrak{m}$-primary integrally closed ideals are Burch ideals.
@@
    If $R$ is of prime characteristic $p$, $F$-finite, and $\operatorname{depth}(R)=1$, then the $e$-th Frobenius pushforward ${{}^{e}\!}{R}$ is Burch for all $e\gg 0$, see [8, 2.4.(vi)]. Recall that ${{}^{e}\!}{R}$ denotes the $R$-module whose underlying abelian group is $R$, and the $R$-action on ${{}^{e}\!}{R}$ is given by $r\cdot x=r^{p^{e}}x$ for $r\in R$ and $x\in{{}^{e}\!}{R}$.
Burch ideals and modules possess several nice homological properties. For instance, they are strong test modules that detect finite homological dimensions via finite vanishing of Ext or Tor ([7, p. 946, Thm. 5.(ii)]; [13, Prop. 3.16]). They characterize regular, complete intersection, Gorenstein, and CM local rings in terms of the corresponding homological dimensions ([7, p. 947, Cor. 1.(ii)]; [13, Cor. 3.19]; [14, Thms. 1.1 and 5.1.(6)]). They also satisfy the Huneke–Wiegand conjecture over Gorenstein rings [8, Thm. 1.2].
Towards understanding growth in minimal free and injective resolutions of Burch modules, it follows from a result of Avramov [6, Thm. 4] that a Burch module (and hence a Burch ideal) has maximal (injective) complexity and curvature; see Proposition 2.20. As a consequence of this and [6, Thm. 3], it follows that $R$ is a complete intersection if (injective) complexity of $M$ is finite, or (injective) curvature of $M$ is at most $1$, where $M$ is some Burch $R$-module. Since $k$ is always a Burch $R$-module, and $\operatorname{cx}_{R}(M)=\operatorname{tcx}_{R}(M,k)=\operatorname{cx}_{R}(M,k)$, $\operatorname{curv}_{R}(M)=\operatorname{tcurv}_{R}(M,k)=\operatorname{curv}_{R}(M,k)$, $\operatorname{inj\,cx}_{R}(M)=\operatorname{cx}_{R}(k,M)$, $\operatorname{inj\,curv}_{R}(M)=\operatorname{curv}_{R}(k,M)$, it is natural to ask the following questions.
@@ 2. Preliminaries
[12, Lem. 2.7 and Rmk. 2.2.(1)] For sequences $\{x_{n}\}$ and $\{y_{n}\}$ of non-negative integers,
@@
The complexity of a module was introduced by Avramov in [4, Def. 3.1] and [5, Def. 1.1], while the injective complexity, referred to as plexity, was defined in [5, Def. 5.1]. The study of (injective) curvature was initiated in [6, Sec. 1].
@@
    [3, p. 286] The complexity (or Ext-complexity) of $M$ and $N$ is defined to be cx_R(M,N):=cx{μ_R(Ext_R^n(M,N))}_n⩾0.
@@
    [12, Def. 3.2] The curvature (or Ext-curvature) of $M$ and $N$ is defined by curv_R(M,N):=curv{μ_R(Ext_R^n(M,N))}_n ⩾0.
@@
    [12, Def. 3.3] The Tor-curvature of $M$ and $N$ is defined by tcurv_R(M,N):=curv{μ_R(Tor^R_n(M,N))}_n ⩾0.
@@
The Tor-curvature is also defined in [19, 3.8] in terms of length of Tor modules. However, the definitions in terms of $\mu$ and in terms of $\lambda$ are equivalent when $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$; see [12, Lem. 3.4].
@@
    [6, Prop. 2] For any $R$-module $M$, the following hold.
@@
    [6, Thm. 3] The following conditions are equivalent:
@@
In view of [18, p. 140, Lem. 2], $\operatorname{Ext}_{R}^{n}(M,k)\cong\operatorname{Ext}_{R/xR}^{n}(M/xM,k)$ and $\operatorname{Ext}_{R}^{n+1}(k,M)\cong\operatorname{Ext}_{R/xR}^{n}(k,M/xM)$ for all $n\geqslant 0$. So $\beta_{n}^{R}(M)=\beta_{n}^{R/xR}(M/xM)$ and $\mu^{n+1}_{R}(M)=\mu^{n}_{R/xR}(M/xM)$ for all $n\geqslant 0$. Hence the desired equalities follow. ∎
@@
(1) It is shown in [3, Thm. II] that $\operatorname{cx}_{R}(M)=\operatorname{inj\,cx}_{R}(M)$ and $\operatorname{cx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}$. By Proposition 2.7.(1), $\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(k)<\infty$. Thus $\operatorname{cx}_{R}(M,N)<\infty$. Hence, in view of Remark 2.6.(3), $\operatorname{curv}_{R}(M,N)\leqslant 1$.
(2) The inequalities for (Tor) complexities are shown in [11, Cor. 5.7] and [12, Prop. 7.9.(1)]. Hence, by Remark 2.6.(3), $\operatorname{tcurv}_{R}(M,N)\leqslant 1$. ∎
@@
(1) Consider the corresponding long exact sequence of Tor modules. Hence, the first claim follows from 2.11, while the desired inequalities can be obtained from 2.11, [12, Lem. 3.4] and Lemma 2.2 by computing the lengths of the Tor modules.
@@
    [13, Def. 3.1] A submodule $M$ of an $R$-module $L$ is said to be Burch if
@@
    [10, Def. 2.1] A Burch submodule of $R$ is called a Burch ideal of $R$.
@@
[13, Lem. 3.6] Let $M$ be a Burch submodule of an $R$-module $L$ such that $\operatorname{depth}(M)>0$. Then there exists an $M$-regular element $x\in\mathfrak{m}$ such that $k$ is a direct summand of $M/xM$.
@@
[6, Thm. 4], [14, Thm. 5.1] Let $M$ be a Burch module. Then
@@
[14, Lem. 4.8] Let $S=R[X]_{\langle\mathfrak{m},X\rangle}$. Let $M$ be a submodule of an $R$-module $L$. Set $L^{\prime}:=L[X]_{\langle\mathfrak{m},X\rangle}$ and $M^{\prime}:=(M+XL[X])_{\langle\mathfrak{m},X\rangle}$. Then
@@
In the common hypothesis of all the statements in [14, Lem. 4.8], it is assumed that $M$ is a Burch submodule of $L$. Note that the Burch property of $M$ is required for the proof of statement (2) in [14, Lem. 4.8], and is not needed elsewhere.
@@
(2) Similarly, as in (1), the desired equalities on injective complexity and curvature can be obtained by using [12, Lem. 3.11.(2)] and Lemma 2.8.(2). ∎
@@
The notion of Gorenstein dimension (in short, G-dimension) is due to Auslander, and it is developed by Auslander and Bridger in [1].
Definition 2.26.
  1. (1)
    An $R$-module $G$ is said to have Gorenstein dimension zero, i.e., $\operatorname{G-dim}_{R}(G)=0$, if $G$ is reflexive, and $\operatorname{Ext}^{i}_{R}(G,R)=\operatorname{Ext}^{i}_{R}(G^{*},R)=0$ for all $i\geqslant 1$, where $G^{*}:=\operatorname{Hom}_{R}(G,R)$.
  2. (2)
    For an $R$-module $M$, $\operatorname{G-dim}_{R}(M)$ is defined to be the infimum of all integers $n\geqslant 0$ such that there exists an exact sequence $0\rightarrow G_{n}\rightarrow G_{n-1}\rightarrow\cdots\rightarrow G_{0}\rightarrow M\rightarrow 0$, where $\operatorname{G-dim}_{R}(G_{i})=0$ for all $0\leqslant i\leqslant n$. If no such $n$ exists, then $\operatorname{G-dim}_{R}(M):=\infty$.
2.27.
Let $\mathcal{A}$ be the category of $R$-modules of Gorenstein dimension zero, and $\mathcal{B}$ be the subcategory of projective $R$-modules. Following the notation as in [2, p. 9], $\hat{\mathcal{A}}$ and $\hat{\mathcal{B}}$ denote the category of $R$-modules of finite Gorenstein dimension and of finite projective dimension, respectively. Then, in view of [2, Thm. 1.1], for each $U$ in $\hat{\mathcal{A}}$, there is an exact sequence $0\rightarrow U\rightarrow V\rightarrow W\rightarrow 0$ with $V$ in $\hat{\mathcal{B}}$ and $W$ in $\mathcal{A}$.